OpenStudy (anonymous):

I have a quick question. to square a number that is written as a power of 10 who do you do to their exponents? can you give an example of how that would be written out. I believe you multiply. But I want to see thanks

OpenStudy (anonymous):

\[ (x^n)^m = \underbrace{x^n\cdot x^n\cdot \ldots x^n}_{m} \]

OpenStudy (anonymous):

\[ =x ^{\overbrace{n+n+\ldots +n}^{m}} = x^{n\cdot m} \]

OpenStudy (anonymous):

I see, I got it I remember this now (10^6 )^2 = 10^12 correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you kindly =)